OpenStudy (anonymous):

There was a thief who went to wine shop to steal wine. Now, when he went, he stole 15 L of wine and added 15 L of water. Again he came and stole 15 L of wine from the same barrel and added 15 L of water. Again he came and stole 15 L of wine from the same barrel and added 15 L of water. At last the ratio of wine to water in the Barrel was 343:169 . Find the initial amount of Wine in the Barrel.

OpenStudy (anonymous):

The answer is supposed to be an integer from 0 to 999 (inclusive). The question seems a little messed up. I assumed that there was no water in the barrel initially. Since 45L wine is removed and 45L of water is added, if we say that the initial amount of wine was 'n', then the ratio at the end of wine to water should be \(\large \frac{n-45}{45}=\frac{343}{169}\). Solving for \( n\) leads to a non-integer value. Which means that the assumption that there is no water initially is incorrect. We now reconsider the problem and assume that there is \(n\) litres of wine initially and \(w\) litres of water initially. At the end when 45L of wine has been added and 45L of water has been removed, the wine to water ratio will be \(\large \frac{n-45}{w+45}=\frac{343}{169}\), which introduces another variable and so we don't enough information to solve the problem...But there is an answer since others have found a correct solution. Any help would be appreciated.

OpenStudy (anonymous):

@satellite73 @primeralph @ganeshie8 @Luigi0210 @myininaya @UnkleRhaukus

ganeshie8 (ganeshie8):

I assumed that there was no water in the barrel initially. is a good assumption.

OpenStudy (anonymous):

you are doing it wrong ....do it step by step, when the thief is stealing wine he gets some water too...so u just cant add or subtract 45 directly

ganeshie8 (ganeshie8):

Since 45L wine is removed and 45L of water is added, Check ocne, 45L wine is NOT removed... when he steals second time, he doesnt get 15L pure wine

OpenStudy (anonymous):

i have done the same question and i was able to solve that...so do it step by step

OpenStudy (anonymous):

OHHHHHHHHHHHHHHHHHHHHHHHH. im shoo shtupid.

OpenStudy (primeralph):

It's just iterative. Nothing difficult.

OpenStudy (anonymous):

lol i see wat u guys mean. the next time he gets a mixture of wine and water and third time as well. wow im dumb i completely forgot lol...i suppose its cus im tired and its almost 2:30 am...now that i know that its a mixture, lets attempt the problem again.

OpenStudy (anonymous):

K I got the answer. It's 120 :) ty for reminding me how stupid i was being at @divu.mkr

OpenStudy (anonymous):

no prob.. (: